substitution method
y=13-3x
2x+9y=-8

Answer by mszlmb(115)

y=13-3x
2x+9y=-8
the first equation tells you that y=13-3x. so, replace y in the second equation with (13-3x).
2x+9(13-3x)=-8.
simple enough. Now solve.[distribute]
2x+117-27x=-8 -117, both sides
-25x=-125 divide by -25
x=5